I think this is proportions yet I am not so sure. 120 decreased by 40% is (__)
Also, 56 INCREASED by 75% is (__)

Answer by jorel1380(3719)

120 decreased by 40% is equal to:
120 (1-40%) = 120(1-.4)=120(.6)=72
56 increased by 75% equals:
56 (1+75%)= 56(1.75) = 98